OpenStudy (anonymous):

When solving x2 – 4x = –7 by completing the square what value is added to both sides of the equation?

OpenStudy (anonymous):

7

OpenStudy (lalaly):

+7

OpenStudy (anonymous):

(b/2)^2 b = -4 -4/2 = -2 -2^2 = 4 4

hero (hero):

pk51 got it. You don't add 7 to both sides to complete the square, you divide b by 2 then square it, then add that result to both sides.